OpenStudy (anonymous):

Evaluate the following (work would be much appreciated): 1. 16^(3/4) 2. (25/4)^(3^2) 3. (4(3)^-2)/((2^-2)(3^-1))

OpenStudy (anonymous):

(2^4 ) ^(3/4) = ...?

OpenStudy (anonymous):

excuse me?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

That's why you convert 16 into 2^4 Now exponent (2^4 ) ^3/4

OpenStudy (anonymous):

umm, im not entirely sure?

OpenStudy (anonymous):

4 * 3/4 = ...?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

thats the answer i got from plugging the original question into a calculator

OpenStudy (anonymous):

i just need the general formula on how to solve those type of problems

OpenStudy (anonymous):

There's no general formula for it, just need practice to get used with exponent!

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

I have to go, but thanks for the help :)

OpenStudy (anonymous):

(2^4 ) ^(3/4) = 2³ = 8